OpenStudy (selena2345):

The ratio of the number Of Taylor's CD To the number of Dave's CDs is 7to8. Taylor has 84CDs. Together How many CDs do they have ?

OpenStudy (anonymous):

The ratio of Taylor's CD to Dave's CD is 7 to 8. If they have a total of 150 CD's. How many CD's does Taylor have? Ratio is 7/8 They share a total of 15 parts of 150 1 part = 150/15 = 10 Taylor has 7 parts of the total He has 7*10=70 CD's AND Dave has 8 partsof the total He has 8*10 = 80 CD's

OpenStudy (anonymous):

k did that help it should have i gave u the answer lol

OpenStudy (selena2345):

yes briannad